# Regularity and temperature of stationary black hole event horizons

### v1

@@ 4 Existence of cross-sections
We stress that in the following Lemma the first instance of “complete” does no refer to affine completeness of some geodesic. Rather it refers to the fact that the parameter $s$ such that $n=\frac{{\rm d}}{{\rm d}s}$ is unbounded from above.
Lemma 4.11.
On a $C^{i}$, $i\geq 3$, null hypersurface $N$ the following two properties are equivalent for a future-directed lightlike tangent vector field $n\in\mathfrak{X}(N)$ which is complete in the future direction.
@@
    $n$ is such that for every $p\in N$ the generator starting at $p$ with tangent $n(p)$ is future incomplete and of finite positive affine length $\Lambda:=-\frac{1}{\kappa}$ where $\kappa<0$ is independent of $p$.
@@
    $n$ satisfies
@@
    for some finite constant $\kappa<0$.
@@
If they hold then the number $\kappa$ mentioned there is the same for both instances and any other $n^{\prime}$ with the same properties satisfies $n^{\prime}=\Lambda^{\prime}/\Lambda n$. Namely, fixing the value of $\Lambda>0$ or $\kappa<0$ fixes the field $n$. A dual statement also holds.
Proof.
Due to (a) the last statement is clear.
(a) $\Rightarrow$ (b). Choose any future-directed lightlike tangent vector field $n^{\prime}$. As its integral curves are the generators, it is necessarily pregeodesic $\nabla_{n^{\prime}}n^{\prime}=\kappa^{\prime}n^{\prime}$, where $\kappa^{\prime}$ is a function. All the different choices for $n^{\prime}$ must necessarily satisfy $01+\kappa^{\prime}\Lambda^{\prime}=\partial_{n^{\prime}}\Lambda=$ 0, see [20, Eq. (12)], thus the existence of $n$ with constant finite $\Lambda>0$, implies that the pregeodesic equation is satisfied with $k=-{1}/{\Lambda}$.
(b) $\Rightarrow$ (a). This is standard, e.g. [20, Lemma 2]. ∎
If we drop the condition that $n$ is complete in the future direction then (a) implies (b) but (b) does not imply (a). This follows again from [20, Eq. (12)].
Definition 4.12.
We say that $N$ has negative surface gravity if $\Lambda$ as in the previous lemma exists. We say that it has positive surface gravity if the previous lemma applies in the time dual version (where negative $\Lambda$ signals past incompleteness). We say that it has zero surface gravity if a geodesic $n$ can be found such that all generators are complete.
Note that the completeness properties are all mutually excluding so the previous definition really defines the sign of $\kappa$.
Remark 4.13.
The point of our definition stands in recognizing the (future/past) completeness condition for $n$, not just in the pregeodesic equation or the affine incompleteness condition. The sign of surface gravity arises from the comparison of two forms of (in)completenesses (in the degenerate case they coincide). The reader can compare our definition with the suggestion by Petersen [41, Def. 1.22] where he imposes the pregeodesic equation and $[k,n]=0$ on $N$. We do not need to impose the latter condition because we shall deduce it, see Thm. 4.14.
@@
In the next result the quotient projection along the generators is still denoted $\tilde{\pi}$, though it applies to $\hat{H}$ rather than $H^{+}$. We recall that $\phi$ is the flow of the Killing field $k$, while $\varphi$ is the flow of a lightlike field $n$ tangent to the horizon.
Theorem 4.14.
@@
Suppose that the event horizon of a stationary black-hole $H^{+}$ is $C^{k}$, $k\geq 3$, totally geodesic and that strong causality holds on $H^{+}$ (or assume the hypothesis of Theorem 3.3) and let the dimension of the spacetime be 4. Let the Euler characteristic of the projection of the event horizon (cf. Def. 3.2) be different from zero, then $\tau>0$ in Cor. 2.19 can be chosen so that $\phi_{\tau}$ sends each generator $\gamma$ to itself (and for any $p\in\gamma$, $\phi_{\tau}(p)>p$), namely on $\hat{H}$ the generators close and $\hat{\pi}:\hat{H}\to S$ is a trivial $S^{1}$-principal bundle over the base manifold $S$ whose fibers are closed Killing orbits.
@@
Moreover, $\tilde{\pi}:\hat{H}\to\tilde{S}$ is a trivial $S^{1}$-principal bundle whose fibers are closed lightlike geodesics. In other words, a future-directed lightlike vector field $n$ tangent to the generators can be found such that its flow $\varphi$ generates an $S^{1}$ action with period $\tau$, $\varphi_{\tau}=Id$ (we say that $n$ is $S^{1}$-generating). Furthermore, after a cycle the tangent to a geodesic generator gets dilated by a factor $e^{c}$ that is independent of the generator (and of the chosen starting point on it).
@@
Additionally, suppose $H^{+}$ is connected, has compact projection (or focus on a connected component with this property) and the spacetime satisfies the dominant energy condition. For every choice of $\tau$ as above, we can find on $\hat{H}$ a $S^{1}$-generating field $n$ with period $\tau$ such that its surface gravity is constant, namely
$\nabla_{n}n=\kappa n,\qquad\kappa=cost,$ (1)
and $c=\kappa\tau$. Moreover, $\kappa$ is uniquely determined.
For $\kappa=0$ all generators are complete, for $\kappa>0$ they are all future complete (and dually).
@@
Let $\check{\tau}$ be the infimum of the possible $\tau$ with the above properties. It is really a minimum (so $\check{\tau}>0$, we speak of rotational case) iff $k$ is not tangent everywhere to the generators, in which case any other $\tau$ is a multiple. Otherwise $\check{\tau}=0$ (irrotational case) and every $\tau>0$ has the above properties. The constant $\kappa$, regarded as a function on $H^{+}$ is uniquely determined as it does not depend on the chosen value of $\tau$. All the vectors fields $n$ for any possible choice of $\tau$ coincide when lifted to $H^{+}$, and if $\check{\tau}=0$, we have $n=k$.
@@
Once $n$ is lifted to $H^{+}$ it is complete and the same equations and properties hold on $H^{+}$. The statement on $n$ being $S^{1}$-generating becomes $\varphi_{\tau}=\phi_{\tau}$ where $\varphi$ and $\phi$ are the flows of $n$ and $k$ respectively.
@@
and if additionally $\tau>0$ the field $\zeta:=\frac{\tau}{2\pi}(k-n)$ provides an $S^{1}$-action on $H^{+}$ with period $2\pi$ (the axisymmetry action). Its non-trivial orbits are spacelike. The flows of the fields $k,n,\zeta$ are isometric in the sense that they preserve the metric induced on the horizon.
@@
If $k$ is not everywhere tangent to the generators, the canonical choice for $\tau$ to realize the compactified space $\hat{H}$ is the minimal possible value $\check{\tau}$. All the physically relevant quantities will be related to this choice and might also be denoted with a ‘check’. The constant $\omega:=2\pi/\tau$ might be called angular velocity but of course the relevant one is the canonical choice, $\check{\omega}:=2\pi/\check{\tau}$.
Our results seem stronger than previous versions, for the control of regularity and for the details of the differential geometric description. Recall that we are not using analyticity, asymptotic conditions, the Einstein’s equations, global hyperbolicity, or the extendibility of $n$ to a Killing field.
Remark 4.15.
The assumption that the spacetime dimension is 4 and the Euler characteristic of the projection is non-zero are really used only to show that $\phi_{\tau}$ closes the generators for some $\tau>0$. Otherwise, the proof is independent of any dimensionality or topological assumptions.
Proof.
The first part of the proof goes precisely as in the proof of [16, Prop. 2.1]. Since $H^{+}$ is totally geodesic the geodesic flow preserves the metric induced on $H^{+}$. Thus on the quotient $\tilde{S}$ we have a well defined Riemannian metric. The Killing vector field $k$ on $(H^{+},g|_{TH^{+}\times TH^{+}})$ (with flow $\phi$) maps generators to generators and descends to a Killing vector field $\tilde{k}$ on $\tilde{S}$, (with flow $\tilde{\phi}$). Observe that $\tilde{k}$ necessarily vanishes at some point, as the Euler characteristic of $\tilde{S}$ is different from zero. Now the argument continues as that given on pp. 119–120 of [47], where it is proved that $\tilde{\phi}_{\tau}=Id_{\tilde{S}}$ for some $\tau>0$ (and a minimal $\tau$ can be chosen if $k$ does not identically project to zero). Consequently, $\phi_{\tau}$ maps each generator to itself.
Since $\hat{H}$ is obtained from $H^{+}$ by quotient with respect to the $\mathbb{Z}$-action generated by $\tilde{\phi}_{\tau}$, the Killing flow induced on $\hat{H}$ generates a $S^{1}\simeq\mathbb{R}/\mathbb{Z}$-free action.
The projection $\hat{\pi}$ is constructed by noticing that every neighborhood of $p\in\hat{H}$ is diffeomorphic with a neighborhood of point in the fiber $c^{-1}(p)$ in the covering space (remember that we have the covering map $c:H^{+}\to\hat{H}$). Which point is chosen is irrelevant as different choices shall have canonically diffeomorphic neighborhoods. Thus composing the local diffeomorphism with $\pi$ we get $\hat{\pi}$ in a neighborhood of $p$. This shows that the bundle $\hat{\pi}:\hat{H}\to S$ does indeed exist.
Similarly, composing the local diffeomorphism with $\tilde{\pi}$ we get that the bundle $\tilde{\pi}:\hat{H}\to\tilde{S}$ does indeed exist. Its fibers are circles. Furthermore, by taking a local section $\Sigma$ passing through $p$, each point of $\Sigma$ is mapped to $\phi_{\tau}(\Sigma)$, a section passing through $\phi_{\tau}(p)$ (each point $q\in\Sigma$ is moved to its causal future, see Thm. 4.1). As $\Sigma$ and $\phi_{\tau}(\Sigma)$ get identified, $\hat{H}$ is covered by products $S^{1}\times\Sigma$, where $\Sigma$ is diffeomorphic to its projection on $\tilde{S}$. In other words, $\tilde{\pi}:\hat{H}\to\tilde{S}$ is a fiber bundle with (oriented) $S^{1}$ fibers. By [39, Prop. 6.15] every oriented $S^{1}$ bundle admits the structure of principal $S^{1}$ bundle. In other words, there is a tangent field $n$ to the generators such that its $n$-parameter over the closed generator has length $\tau$ (regardless of the generator).
Let $n$ be a $S^{1}$-generating field with period $\tau$. Since the bundle $\tilde{\pi}:\hat{H}\to\tilde{S}$ is trivial we can introduce a section and push it with the $S^{1}$-isometry to get a foliation. Call the tangents spaces to this foliation horizontal. They are preserved by the flow $\varphi$ of $n$. We can introduce a 1-form field $n^{*}$ such that $n^{*}(n)=1$ and its kernel is horizontal. At this point we proceed with the ribbon argument, that is, let us consider two integral curves $x_{0}(s)$ and $x_{1}(s)$ of $n$ starting from one level set $\Sigma$ and terminating in another level set $\Sigma^{\prime}$, we can close the starting points with a horizontal curve and the ending points with another horizontal curve (which is the image under the flow of the former), and we have by [20, Eq. (18)], by choosing the longitudinal length to be that of $m$ cycles
@@
This integral has the following interpretation. Let us consider the affinely parametrized geodesic $\gamma(t)$ with initial conditions $\gamma(0)=x(0)$ and $\dot{\gamma}(0)=n(x(0))$. We have $\dot{\gamma}(x(s))=f(s)n(x(s))$, and from the geodesic condition $\nabla_{\dot{\gamma}}\dot{\gamma}=0$ it follows [20, proof of Lemma 2]

### v2

@@ 6 The (sign of) surface gravity (including the non-compact case)
In the next result “forward complete” does not refer to the affine completeness of some geodesic. Rather it refers to the fact that the parameter $s$ such that $n=\frac{{\rm d}}{{\rm d}s}$ over the geodesic is unbounded from above. For a previous result in the direction of the following theorem, see (costa24, , Prop. 3.3).
Theorem 6.1
On a $C^{0}$ future null hypersurface $N$ the following two properties are equivalent for a regular semi-tangent section $n$
@@
    there is a finite constant $\kappa<0$ such that for every $p\in N$ any generator starting at $p$ with tangent belonging to the semi-tangent section has finite affine length $\Lambda=-\frac{1}{\kappa}$ in the future direction,
@@
    there is a finite constant $\kappa<0$ such that over every generator of NN
@@
    and $n$ is forward complete.
@@
If they hold then the number $\kappa$ mentioned there is the same for both instances and any other regular semi-tangent section $n^{\prime}$ with the same properties satisfies $n^{\prime}=\frac{\Lambda}{\Lambda^{\prime}}\,n$. Namely, fixing the value of $\Lambda>0$ or $\kappa<0$ fixes the field $n$. Conversely, fixing the scale for $n$ fixes $\kappa$ and $\Lambda$. A dual statement also holds.
Proof
Let us focus on a generator and let us denote with $n$ the $C^{1}$ tangent field over it coming from the regular semi-tangent section. As shown above $\nabla_{n}n=\alpha n$, where $\alpha$ is a function. Let $x:[0,b)\to N$ , $s\mapsto x(s)$, be the (generator) integral curve of $n$, with $x(0)=p$. As proved in (minguzzi21, , Lemma 2) (for the case $b=+\infty$ but the proof does not change), the geodesic $\gamma$ starting from $x(\tau)$ with tangent $\dot{\gamma}=n$ has future affine length
$\Lambda(x(\tau))=\int_{\tau}^{b}e^{\int_{\tau}^{\rho}\alpha(x(s)){\rm d}s}{\rm d}\rho.$ (1)
In particular, $1+\alpha\Lambda+n(\Lambda)=0$.
(a) $\Rightarrow$ (b). Evidently, the assumption (a) implies that $\Lambda$ is constant over the generator and so that the last term in the left-hand side of $1+\alpha\Lambda+n(\Lambda)=0$ vanishes. This implies that $\kappa=\alpha$. From (1) we get as, $\alpha$ is constant,
$\Lambda(x(\tau))=\int_{\tau}^{b}e^{\kappa(\rho-\tau)}{\rm d}\rho=\frac{e^{\kappa(b-\tau)}-1}{k}.$ (2)
Suppose that $b<+\infty$, then $\Lambda(x(\tau))$ depends on $\tau$ and so $\Lambda$ is not constant. The contradiction proves that $b=+\infty$ and hence that $n$ is forward complete. Since $\alpha=\kappa$ we also get $\nabla_{n}n=\alpha n=\kappa n$.
(b) $\Rightarrow$ (a). If $b=+\infty$ and $\nabla_{n}n=\kappa n$ then $\alpha=\kappa$ thus again integrating (1) we get $\Lambda=-1/\kappa$ which does not depend on the starting point on $N$ and on the generator chosen.
Due to (a) the last statement is clear. $\square$
We arrive at the definition of surface gravity. Note that we are not interested in assigning a number to each point of $N$ (this is known as dynamical surface gravity among physicists pielahn11 ) but rather a number to $N$.
Definition 3 (Surface gravity $\kappa$ and its sign)
We say that a $C^{0}$ future null hypersurface $N$ has negative surface gravity if there exists a semi-tangent section $n$ with the properties of Theorem 6.1. We say that a $C^{0}$ past null hypersurface $N$ has positive surface gravity if Theorem 6.1 applies in the time dual version. We say that a $C^{1}$ null hypersurface has zero surface gravity if a geodesic field $n$ can be found such that all generators are complete. When the scale of $n$ can be fixed the surface gravity $\kappa$ is defined. It is zero if the last case apply, otherwise it can be read from $\nabla_{n}n=\kappa n$ or $\kappa=-1/\Lambda$.
Note that a $C^{0}$ future null hypersurface is also a $C^{0}$ past null hypersurface iff it is a $C^{1}$ null hypersurface. In fact, by local achronality every future generator starting at $p$ must match (no corner) a past generator at $p$, which implies that every point is in the interior of a generator.
Even in the $C^{1}$ case the affine completeness properties are all mutually excluding so the previous definition is consistent. The formula $\Lambda=-1/\kappa$ holds also in the positive surface gravity case, where a negative $\Lambda$ signals past incompleteness.
Note that if $n$ is geodesic then completeness is equivalent to affine completeness.
In the following section we shall show how the scale of $n$ is fixed for the event horizons.
Remark 3
Theorem 6.1 allows us to recognize the importance of the completeness of the field $n$ for the definition of the sign of surface gravity. It is not sufficient to refer to the pregeodesic equation $\nabla_{n}n=\kappa n$ with constant $\kappa$ to read the value of the sign surface gravity, for it is necessary to check that $n$ is forward complete for $\kappa<0$ (an inequality which implies future affine incompleteness), backward complete for $k>0$ (an inequality which implies past affine incompleteness), and complete for $\kappa=0$. This check is not needed for compact $N$, as the completeness property would be ensured by standard ODE theory, but is otherwise necessary.
For instance, over a smooth event horizon which admits a section we can find a geodesic $n$. It is sufficient to fix the value of $n$ on the section and then propagate it geodetically. Though in this way we get $\nabla_{n}n=0$ we cannot conclude that $\kappa=0$ as $n$ might be non-complete.
Therefore, the sign of surface gravity is related to two contrasting forms of (in)completenesses (which, actually, in the degenerate case coincide). The reader can compare our definition with the suggestion by Petersen (petersen19, , Def. 1.22) where the pregeodesic equation is paired instead with a condition of the form, $[k,n]=0$ on $N$, with $k$ Killing. We shall meet this type of property later on but it will be deduced, see Thm. 7.2.
7 The surface gravity (temperature) and angular velocity of event horizons
The goal of this section is to explore in some detail the geometry of the horizon in 4-dimensional spacetimes under the dominant energy condition. In any case, all assumptions will be spelled out in the theorems below. Some results and strategies can actually work in other dimensions.
Our main result establishes that it is possible to assign a constant surface gravity (not just the sign), to be interpreted as temperature via the formula $T=\frac{\kappa}{2\pi}$, an angular velocity, a privileged lightlike vector field $n$ tangent to the horizon, and a privileged vector field $\zeta$ on the horizon generating an $S^{1}$-action (axisymmetry). These fields are related as follows:
$k=n+\omega\zeta.$
These results are not entirely new, of course, the statement concerning the existence of an axisymmetry being related to what in the literature goes under the name of black hole rigidity theorem.
It is worth recalling the general traditional strategy for this type of investigation, also to emphasize the differences from our findings. The standard approach, due to Hawking hawking72 ; chrusciel97 (hawking73, , Prop. 9.3.6), is to prove under the following conditions:
  • (a)
    analyticity of the metric and horizon,
  • (b)
    (electro-)vacuum assumption or assumptions on the hyperbolic evolution of matter from Cauchy data,
  • (c)
    asymptotic flatness conditions and causality assumptions,
  • (d)
    energy conditions,
  • (e)
    bifurcation conditions,
that a stationary spacetime admitting a Killing field $k$ which is timelike near the conformal boundary must have a Killing event horizon and must be axisymmetric (the Killing field $n$ tangent to the horizon might be different from $k$). The $S^{1}$-action Killing field $\zeta$ is first found in a neighborhood of the horizon and then extended by using real analyticity. The Killing property of the horizon also implies constant surface gravity, via simple algebraic computations making use of the dominant energy condition bardeen73 ; chrusciel20 . Similar rigidity results hold in higher dimensions hollands07 ; hollands09 .
Once these results are established, other arguments, not depending on the analyticity assumption, first explored by Carter carter71b and Robinson robinson75 , show that the black hole is determined by just three parameters: mass, angular momentum, and electric charge. This result is known as the no-hair theorem. This line of inquiry thus leads to the proof of the black hole uniqueness theorem in the analytic case, which essentially states that the only spacetime with the above (suitable) properties is the Kerr(-Newman) solution robinson75 ; mazur82 ; heusler96 ; amsel10 ; chrusciel10b ; meinel12 ; chrusciel12b ; chrusciel20 . Incidentally, by our Theorem 3.2, any analytical regularity assumption on the horizon, as mentioned in (a), can be dropped.
Our main contribution in this section goes in a different direction. To start with, we do not impose an electro-vacuum assumption or assumptions on the hyperbolic evolution of matter from Cauchy data. Even asymptotically, we do not require the spacetime to be vacuum, just stationary. We can imagine a stationary configuration in which an ingoing matter-energy flow balances a similar outflow. There can still be an event horizon, in which case our result shows that we can assign a temperature and an angular velocity to it.
We do not impose any bifurcation assumption, thus completely dropping (a), (b), and (e). Of course, our results are correspondingly weaker. Indeed, we stress that the philosophy of our approach is different from Hawking’s. We do not try to prove that the horizon is Killing from the outset; in fact, such a problem is difficult in the smooth case moncrief83 ; isenberg85 ; friedrich99 ; racz00 ; ionescu09 ; alexakis10 ; petersen19 , and has been solved so far only in the non-degenerate case and again under vacuum assumptions (see the discussion at the end of the section; in any case, typically one needs to prove at the very first step that the surface gravity is constant without using the analyticity assumption). Instead, we study the geometry of the horizon in itself and show the existence of the isometries just on the horizon. The temperature and angular velocity are thus shown to be intrinsic to the horizon. [^1] In a way, our approach is closer in spirit to the study of near-horizon geometry kunduri13 ; dunajski25 , for in that setting one focuses on the intrinsic geometry of the horizon as it follows from the embedding in the spacetime by somehow projecting/reducing the Einstein equations to it (but in those works the horizon is assumed to be Killing and degenerate from the outset, and the vacuum equations are imposed, all assumptions that we do not impose).
If we assume an (electro-)vacuum condition, our findings are connected to the attempt to prove the black hole uniqueness theorem in the smooth framework. As is well known, there has been only limited success as analyticity cannot be used to extend various objects obtained at and near the horizon. Suppose such uniqueness is false. Our result shows that, nevertheless, we can still assign a temperature and an angular velocity to the horizon regardless of it being Killing.
Of course, there is another take on our results. Our hope is that they might ultimately help to prove the black hole uniqueness theorem in the smooth setting. The idea is to start constraining the geometry of the horizon, and learning as much as possible about it, as we try to do in this section. From here, PDE techniques could hopefully be used to get the relevant extension over the whole domain of outer communication.
We shall make use of the following result.
Lemma 8
Let $(N,g)$ be a compact Riemannian manifold, $n$ a Killing field such that $g(n,n)=1$, and $k$ a vector field such that $[k,n]=fn$ for some function $f:N\to\mathbb{R}$ such that $\partial_{n}f=0$. Then $f=0$ and hence $[k,n]=0$.
Proof
We denote with $\nabla$ the Levi-Civita connection of $g$. From $[k,n]=fn$, we have, taking the scalar product with $n$
${g}({\nabla}_{k}n,n)-{g}({\nabla}_{n}k,n)=f.$
As $n$ is a unit Killing vector field for (N,g)(N,{g}), ${\nabla}_{n}n=0$, so we get n(g(k,n))=f-n({g}(k,n))=f. From $[k,n]=fn$, we also have, taking the scalar product with $k$
${g}({\nabla}_{k}n,k)-{g}({\nabla}_{n}k,k)=f{g}(k,n).$
Since $n$ is a Killing vector field for (N,g)({N},{g}), we get n(g(k,k))=2fg(k,n).n({g}(k,k))=-2f{g}(k,n). From this we get n(n(g(k,k)))=2fn(g(k,n))n(n({g}(k,k)))=-2fn({g}(k,n)) since $n(f)=0$. But n(g(k,n))=f-n({g}(k,n))=f, so it follows that $n(n({g}(k,k)))=2f^{2}$. Let us set $h:={g}(k,k):N\to\mathbb{R}$ which, being continuous and defined on a compact set, is bounded. For any integral curve $\gamma:\mathbb{R}\to N$ of $n$, we have $(h\circ\gamma)^{\prime\prime}(t)\geq 0,\forall\,t\in\mathbb{R};$ hence $h\circ\gamma$ is a convex function defined on $\mathbb{R}$, but since $h\circ\gamma$ is bounded, it is constant. It follows that for any integral curve $\gamma$ of $n$, $(h\circ\gamma)^{\prime\prime}(t)=0,\forall\,t\in\mathbb{R}$; which proves that $f=0$. Hence $[k,n]=0$. $\square$
@@ 7 The surface gravity (temperature) and angular velocity of event horizons
In the next result the quotient projection along the generators is always denoted $\tilde{\pi}$, even when it applies to the compactified horizon $\hat{H}$ rather than $H^{+}$. We recall that $\phi$ is the flow of the Killing field $k$, while $\varphi$ is the flow of a lightlike field $n$ tangent to the horizon.
For simplicity, we formulate the following two theorems for a connected event horizon $H^{+}$, though they hold equally well for a connected component of $H^{+}$, provided it satisfies the properties mentioned in the theorems.
Theorem 7.1 (Compactified horizon with closed generators)
Let $(M,g)$ be a 4-dimensional spacetime endowed with a $C^{k}$, $k\geq 3$, metric. Let $k$ be a complete Killing vector field. Let $\Sigma_{end}$ be an acausal hypersurface, possibly with edge, over which $k$ is timelike. Let $M_{end}:=\phi(\Sigma_{end})$ and suppose that the horizon $H^{+}:=\partial I^{-}(M_{end})\cap I^{+}(M_{end})$, is connected, has compact Killing projection $S=\pi(H^{+})$ (cf. Def. 2) of non-vanishing Euler characteristic, and that strong causality holds on $H^{+}$. Furthermore, assume that $H^{+}$ is as regular as the metric and totally geodesic (or assume $\theta\geq 0$ in support sense over $H^{+}$ and the null convergence condition so as to apply Theorem 3.2).
@@
Then the constant $\tau>0$ in Cor. 5 can be chosen so that $\phi_{\tau}$ sends each generator $\gamma$ to itself (and for any $p\in\gamma$, $\phi_{\tau}(p)>p$), namely on $\hat{H}$ the generators close and $\hat{\pi}:\hat{H}\to S$ is a trivial $S^{1}$-principal bundle over the base manifold $S$ whose fibers are closed Killing orbits.
@@
Moreover, $\tilde{\pi}:\hat{H}\to\tilde{S}$ is a trivial $S^{1}$-principal bundle whose fibers are closed lightlike geodesics. More precisely, a future-directed lightlike vector field $n$ tangent to the generators can be found such that its flow $\varphi$ generates an $S^{1}$-action with period $\tau$, $\varphi_{\tau}=Id$ (we say that $n$ is $S^{1}$-generating).
@@
Let $\check{\tau}$ be the infimum of the possible $\tau$ with the above properties. It is really a minimum (so $\check{\tau}>0$, and we speak of rotational case) iff $k$ is not tangent everywhere to the generators, in which case any other $\tau$ is a multiple. Otherwise $\check{\tau}=0$ (irrotational case) and every $\tau>0$ has the above properties.
@@
The vector field $n$ lifts from $\hat{H}$ to $H^{+}$ to a complete vector field (denoted in the same way). The statement on $n$ being $S^{1}$-generating on $\hat{H}$ with period $\tau$ becomes on $H^{+}$, $\varphi_{\tau}=\phi_{\tau}$, where $\varphi$ and $\phi$ are the flows of $n$ and $k$ respectively.
Proof
The first part of the proof goes precisely as in the proof of (friedrich99, , Prop. 2.1). Since $H^{+}$ is totally geodesic the geodesic flow preserves the metric induced on $H^{+}$. Thus on the quotient $\tilde{S}$ we have a well defined Riemannian metric. The Killing vector field $k$ on $(H^{+},g|_{TH^{+}\times TH^{+}})$ (with flow $\phi$) maps generators to generators and descends to a Killing vector field $\tilde{k}$ on $\tilde{S}$, (with flow $\tilde{\phi}$). Observe that $\tilde{k}$ necessarily vanishes at some point, as the Euler characteristic of $\tilde{S}$ is different from zero. The next step in the argument is as in pages 119–120 of wald94 , where it is proved that $\tilde{\phi}_{\tau}=Id_{\tilde{S}}$ for some $\tau>0$ (and a minimal $\tau$ can be chosen if $k$ does not identically project to zero, in which case $k$ and $n$ are proportional on $H^{+}$). Consequently, $\phi_{\tau}$ maps each generator to itself.
Since $\hat{H}$ is obtained from $H^{+}$ by quotient with respect to the $\mathbb{Z}$-action generated by $\phi_{\tau}$, the Killing flow induced on $\hat{H}$ generates a $S^{1}\simeq\mathbb{R}/\mathbb{Z}$-free action.
The projection $\hat{\pi}$ is constructed by noticing that every sufficiently small neighborhood of $p\in\hat{H}$ is diffeomorphic with a neighborhood of point in the fiber $c^{-1}(p)$ in the covering space (remember that we have the covering map $c:H^{+}\to\hat{H}$). Which point is chosen is irrelevant as different choices shall have canonically diffeomorphic neighborhoods. Thus composing the local diffeomorphism with the projection $\pi$ constructed in Cor. 3 we get $\hat{\pi}$ in a neighborhood of $p$. This shows that the bundle $\hat{\pi}:\hat{H}\to S$ does indeed exist.
Similarly, composing the local diffeomorphism with $\tilde{\pi}$ and using Thm. 5.1 we get that the bundle $\tilde{\pi}:\hat{H}\to\tilde{S}$ does indeed exist. Its fibers are circles. Furthermore, by taking a local section $\Sigma$ of $\tilde{\pi}:H^{+}\to\tilde{S}$ passing through $p$, each point of $\Sigma$ is mapped to $\phi_{\tau}(\Sigma)$, a section passing through $\phi_{\tau}(p)$ (each point $q\in\Sigma$ is moved to its causal future, see Thm. 4.1). As $\Sigma$ and $\phi_{\tau}(\Sigma)$ get identified, $\hat{H}$ is covered by products $S^{1}\times\Sigma$, where $\Sigma$ is diffeomorphic to its projection on $\tilde{S}$. In other words, $\tilde{\pi}:\hat{H}\to\tilde{S}$ is a fiber bundle with (oriented) $S^{1}$ fibers. By (morita01, , Prop. 6.15) every oriented $S^{1}$ bundle admits the structure of principal $S^{1}$ bundle. In other words, there is a tangent field $n$ to the generators such that its $n$-parameter over the closed generator has length $\tau$ (regardless of the generator).
[…]
If $k$ is not everywhere tangent to the generators, the canonical choice for $\tau$ to realize the compactified space $\hat{H}$ is the minimal possible value $\check{\tau}$. All the physically relevant quantities will be related to this choice and might also be denoted with a ‘check’. We call the constant $\check{\tau}$ the rotational period of the horizon ad the constant $\check{\omega}:=2\pi/\check{\tau}$ the angular velocity of the horizon. The latter is related to a $S^{1}$-action on $H^{+}$ as clarified by the following theorem.
Working with $\tau$ might be convenient to avoid dealing with the cases $\check{\tau}>0$ and $\check{\tau}=0$ separately.
Note that $n$ is complete in $H^{+}$, thus in $H$ it cannot be extended to a vector field unless it converges to zero on $\textrm{edge}(H^{+})$.
The proof of the following result uses the compactification of Theorem 7.1, however, for better readability, the statement does not mention such construction.
Theorem 7.2 (Lightlike field, surface gravity and angular velocity of the horizon)
Assume the hypothesis of Theorem 7.1 and suppose, furthermore, that the dominant energy condition holds.
On the horizon $H^{+}$ there is one and only one future-directed $C^{k-1}$ lightlike field $n$, called the lightlike field of the horizon, with the properties:
  • (a)
    completeness,
  • (b)
    pregeodesic equation
    $\nabla_{n}n=\kappa n,$ (3)
    for some constant $\kappa$,
  • (c)
    $\varphi_{\tau}=\phi_{\tau}$ ($S^{1}$-generating condition) for some $\tau>0$.
The constant $\kappa$ for the lightlike field of the horizon is thus uniquely determined and called the surface gravity of the horizon.
For $\kappa=0$ all generators are (affine) complete. For $\kappa>0$ they are all future affine complete, and past affine incomplete in the portion contained in $H^{+}$ (and time dually for $\kappa<0$).
Further, there are two cases. Either $n=k$ in which case (c) holds for any $\tau$, or $n\neq k$ somewhere on $H^{+}$ in which case (c) holds for $n$ for a minimal value of $\tau$ denoted $\check{\tau}>0$. In the former case we set $\check{\omega}:=0$ and in the latter case we set $\check{\omega}:=2\pi/\check{\tau}$, and call $\check{\omega}$ the angular velocity of the horizon.
Consider a lightlike geodesic $\gamma:I\to H^{+}$ running on the horizon starting with velocity $n$. Then we have the vector ratio $\dot{\gamma}(\check{\tau})/d\phi_{\check{\tau}}(\dot{\gamma}(0))=e^{-\check{c}}$ where the constant is $\check{c}=\kappa\check{\tau}$ and so is independent of the starting point (note also that $d\phi_{\check{\tau}}(n)=d\varphi_{\check{\tau}}(n)=n$). We call $e^{-\check{c}}$ the dilating factor of the horizon.
@@
and if $\check{\tau}>0$ the field $\zeta:=\frac{\check{\tau}}{2\pi}(k-n)$ provides a $S^{1}$-action on $H^{+}$ with period $2\pi$ (the axisymmetry action; we can also set $\zeta:=0$ for $\check{\tau}=0$). Its non-trivial orbits are spacelike. The flows of the fields $k,n,\zeta$ are isometric in the sense that they preserve the metric induced on the horizon.
@@
The proof is somewhat long so we organize it in sections with titles in boldface.
Proof
We use the compactification $\hat{H}$ relative to the parameter $\tau>0$ introduced in Theorem 7.1.
The dilating factor and its generator independence. Let $n$ be a $S^{1}$-generating field with period $\tau$ on $\hat{H}$. Since the bundle $\tilde{\pi}:\hat{H}\to\tilde{S}$ is trivial we can introduce a section $\Sigma$ and push it with the $n$-flow $\varphi$ to get a foliation. Call the tangents spaces to this foliation horizontal. They are preserved by the flow $\varphi$ of $n$. We can introduce a 1-form field $n^{*}$ such that $n^{*}(n)=1$ and its kernel is horizontal. At this point we proceed with the ribbon argument, that is, let us consider two integral curves $x_{0}(s)$ and $x_{1}(s)$ of $n$ starting from $\Sigma$ and terminating in another leaf $\Sigma^{\prime}$. We can connect the starting points with a horizontal curve in $\Sigma$ and the ending points with another horizontal curve in $\Sigma^{\prime}$ which is the image under the flow of the former (note that since $H^{+}$ is connected, $\tilde{S}$ is connected), and we have by (minguzzi21, , Eq. (18)) (which uses the dominant energy condition), by choosing the longitudinal length to be that of $m$ cycles
@@
This integral has the following interpretation. Let us consider the affinely parametrized geodesic $\gamma(t)$ with initial conditions $\gamma(0)=x(0)$ and $\dot{\gamma}(0)=bn(x(0))$ for some $b>0$. We have $\dot{\gamma}(t(s))=f(s)bn(x(s))$, and from the geodesic condition $\nabla_{\dot{\gamma}}\dot{\gamma}=0$ it follows (minguzzi21, , proof of Lemma 2)